# Connecting the Dots: A Chain-of-Collaboration Prompting Framework for LLM Agents

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive performance in executing complex reasoning tasks. Chain-of-thought effectively enhances reasoning capabilities by unlocking the potential of large models, while multi-agent systems provide more comprehensive solutions by integrating collective intelligence of multiple agents. However, both approaches face significant limitations. Single-agent with chain-of-thought, due to the inherent complexity of designing cross-domain prompts, faces collaboration challenges. Meanwhile, multi-agent systems consume substantial tokens and inevitably dilute the primary problem, which is particularly problematic in business workflow tasks. To address these challenges, we propose Cochain, a collaboration prompting framework that effectively solves business workflow collaboration problem by combining knowledge and prompts at a reduced cost. Specifically, we construct an integrated knowledge graph that incorporates knowledge from multiple stages. Furthermore, by maintaining and retrieving a prompts tree, we can obtain prompt information relevant to other stages of the business workflow. We perform extensive evaluations of Cochain across multiple datasets, demonstrating that Cochain outperforms all baselines in both prompt engineering and multi-agent LLMs. Additionally, expert evaluation results indicate that the use of a small model in combination with Cochain outperforms GPT-4.
@@
1 Instruction
Large Language Models (LLMs) have demonstrated impressive performance in language understanding and complex reasoning tasks [51, 5, 21, 8], with prompt engineering playing a crucial role in enhancing their reasoning capabilities. Prompt engineering represented by Chain-of-Thought prompt LLMs in ways that facilitate coherent and step-by-step reasoning processes, endowing agents with depth of thought [7, 57, 66]. To enhance agent collaborative capabilities, various multi-agent collaborative frameworks have been developed, providing agents with breadth of thought. These frameworks encode carefully designed agent profiles and collaborative mechanisms into prompts, yielding favorable outcomes across domains such as healthcare [50, 29, 38], education [17, 44, 22], law [16], and finance [64]. Due to their exceptional problem-solving and collaborative abilities, multi-agent systems are considered a promising pathway toward Artificial General Intelligence (AGI) [9, 81]. However, this raises a critical question: Is a multi-agent LLM system all you need?
Evidently not, as both single-agent and multi-agent systems exhibit limitations [53]. As illustrated in Figure 1, the inherent complexity in designing cross-domain prompts [47] means that single agents employing prompt engineering lack constraint awareness and face challenges in cross-domain collaboration (under-collaboration). Simultaneously, multi-agent systems face challenges from expensive token consumption and extensive inference time [20, 60], with inter-agent communications often ineffectively utilized [72]. More importantly, existing research primarily focuses on maximizing collaboration, extensively exploring how multi-agent systems can improve decision quality [73, 43], make safe decisions [33, 42], and solve problems [78, 34], while paying insufficient attention to the negative effects of excessive agent collaboration. When all agents simultaneously participate in every decision-making process, core issues are frequently overwhelmed by non-critical information, leading not only to a decline in answer quality but also to a preference for consensus-driven answers over accurate ones [20], particularly in business workflow tasks. To our best knowledge, no existing literature has systematically analyzed and defined this “over-collaboration” phenomenon or proposed targeted solutions. This prompts us to consider: Why not combine the reasoning capabilities of chain-of-thought prompting with the collaborative nature of multi-agent systems?
Inspired by these challenges, we propose a collaborative prompting framework, Cochain, designed to enhance LLM agents’ collaborative capabilities in business workflow tasks at low cost, while mitigating over-collaboration through indirect rather than direct agent participation in decision-making. As shown in Figure 2, Cochain comprises two primary components: collaborative knowledge graph and prompts tree. We mine agent knowledge through counterfactual reasoning and refine it into a knowledge graph integrated with explicit dataset knowledge. To mitigate the impact of irrelevant knowledge on model performance [79], we introduce causal chain structure to improve cross-stage reasoning capabilities. Inspired by how the human brain integrates fragmented knowledge into coherent thought [15], we propose prompts tree that distills prompts from agent responses. By retrieving from the prompts tree, agents can automatically acquire chain-like cross-stage prompts, avoiding the high cost of manual prompt design while effectively utilizing inter-agent communications.
Figure 1: The impact of agent collaboration degree on problem-solving effectiveness. Under-collaboration ignores constraints from other stages, while over-collaboration dilutes main issues with irrelevant information (underlined in our examples). Agents using Cochain achieve balance by considering constraints while maintaining focus on key problems, effectively connecting the dots.
@@ 1 Instruction
  • $\bullet$
    Effective Collaboration: We propose an agent collaborative prompting framework, Cochain, that equips agents with both depth and breadth of thinking, effectively alleviating the problems of over-collaboration and under-collaboration.
  • $\bullet$
    Cost-Effective yet Efficient: Our extensive experimental results demonstrate that Cochain exhibits significant improvements over all baselines across different model backbones and three business workflow datasets. Cochain possesses reasoning speeds comparable to single-agent systems and even lower inference costs, and surpasses existing multi-agent systems in capability, achieving an optimal balance between performance and efficiency.
  • $\bullet$
    The Small + Cochain Outperforms the Large: Evaluations by domain experts have further confirmed that the combination of the small model and our framework outperforms the large model. To support further research on over-collaboration and business workflow tasks, we will publicly release our collaborative knowledge graph and high-quality datasets.
@@ 2 Related Work
Figure 2: Cochain connects business workflow stages. In part (a), agents address counterfactual inquiries, with answers undergoing interactive iteration. In part (b), a prompts tree is built. Part (d) presents our framework: an agent uses the prompts tree and a collaborative knowledge graph built from counterfactual answers and original data, achieving more collaborative results than part (c).
@@ 2.1 Knowledge Graph Augmented LLMs
LLM reasoning capabilities are critical for high-quality responses [28, 49, 30]. Knowledge graphs, with their structured, explicit, and interpretable nature [75, 40, 80], offer clear representations and transparent reasoning paths, effectively guiding LLMs in deeper reasoning and mitigating limitations [36, 67, 56]. For domain-specific LLMs, knowledge graphs enable continuous updating of knowledge frameworks to accommodate evolving information [27, 37, 31]. Currently, MindMap [59] applies knowledge graphs in medicine, while ChatLAW [16] employs them in legal domains for process simulation. Business workflows, due to their multi-stage nature and cross-domain integration requirements, face the unique challenge of multi-stage knowledge fusion. To prevent performance degradation from irrelevant knowledge [39], we propose a causal chain optimizing knowledge selection, enhancing understanding of cross-stage relationships while filtering out noise.
@@ 2.2 Prompt Engineering
Prompt engineering aims to guide LLMs to maximize their potential through carefully designed prompts [11, 47, 76]. Its core enables models to adapt to new tasks by leveraging contextual learning and following instructions [10, 32]. Research shows well-designed prompts significantly enhance model performance across tasks, particularly complex ones [57, 55, 77, 65]. CoT [57] enables models to generate intermediate reasoning steps through simple prompts, improving performance. ToT [66] explores coherent text units as intermediate problem-solving steps. CPO [71] leverages non-optimal reasoning paths generated through tree search, speeding up inference. However, existing prompt engineering methods lack sustained constraint awareness and neglect cross-domain collaboration. Cochain enables LLM agents to address this challenge while balancing interdependent constraints and implementation feasibility.

### v2

# Cochain: Balancing Insufficient and Excessive Collaboration in LLM Agent Workflows

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive performance in executing complex reasoning tasks. Chain-of-thought effectively enhances reasoning capabilities by unlocking the potential of large models, while multi-agent systems provide more comprehensive solutions by integrating the collective intelligence of multiple agents. However, both approaches face significant limitations. Single-agent with chain-of-thought, due to the inherent complexity of designing cross-domain prompts, faces collaboration challenges. Meanwhile, multi-agent systems consume substantial tokens and inevitably dilute the primary problem, which is particularly problematic in business workflow tasks. To address these challenges, we propose Cochain, a collaboration prompting framework that effectively solves the business workflow collaboration problem by combining knowledge and prompts at a reduced cost. Specifically, we construct an integrated knowledge graph that incorporates knowledge from multiple stages. Furthermore, by maintaining and retrieving a prompts tree, we can obtain prompt information relevant to other stages of the business workflow. We perform extensive evaluations of Cochain across multiple datasets, demonstrating that Cochain outperforms all baselines in both prompt engineering and multi-agent LLMs. Additionally, expert evaluation results indicate that the use of a small model in combination with Cochain outperforms GPT-4.
@@
1 Introduction
Figure 1: Collaboration has a Goldilocks region. With Insufficient Collaboration, a system overlooks cross-stage constraints and dependencies. With Excessive Collaboration, interaction injects peripheral details that dilute the core decision and increase token cost. Cochain targets effective collaboration by enforcing stage constraints while keeping the solution focused. We illustrate this with a car manufacturing query, where a single agent misses manufacturing and supplier constraints, and excessive discussion dilutes focus on issues such as inventory rollout and price volatility.
Large Language Models (LLMs) have achieved strong performance on language understanding and complex reasoning tasks Touvron et al. (2023a); Bai et al. (2023); GLM et al. (2024); Bi et al. (2024). Much of this progress is unlocked at inference time. Prompting techniques such as Chain-of-Thought encourage step-by-step reasoning and often improve depth of reasoning Wei et al. (2022); Besta et al. (2024). In parallel, multi-agent frameworks extend a single model into a team of specialized roles and interaction protocols, aiming to provide broader coverage across domains such as healthcare, education, law, and finance Tang et al. (2024); Dan et al. (2023); Cui et al. (2024); Yang et al. (2023). However, a practical question remains: How much collaboration is actually needed for high-quality decision making?
This question is especially salient in business workflow tasks. We define a Business Workflow as a complex process consisting of many interconnected stages. The stages are specialized, and they strongly depend on each other. A decision made in one stage requires more than local expertise. It must anticipate and incorporate knowledge and constraints from other stages. For example, in an automotive manufacturing workflow, design and production choices must align with what the supply chain can actually provide. In this setting, we observe two opposite failure modes (Figure 1). Insufficient Collaboration arises when the overall system does not effectively integrate downstream constraints into the decision process, so solutions may look locally reasonable but violate workflow level requirements Sahoo et al. (2024). Excessive Collaboration arises when many agents participate directly and simultaneously in each decision, making interaction expensive and unfocused, so non critical information overwhelms the main issue and the system may prefer consensus-like answers over accurate ones Du et al. (2024); Zhang et al. (2024b). Existing research on multi-agent systems primarily focuses on maximizing collaboration, extensively exploring how multi-agent systems improve decision quality Zhang et al. (2025b), make safe decisions Piatti et al. (2024), and solve complex problems Li et al. (2025). In contrast, the negative effects of excessive collaboration have received insufficient attention, particularly in multi-stage business workflows Lei et al. (2025). To our best knowledge, no prior work has systematically analyzed and defined this Excessive Collaboration phenomenon in workflow settings, nor proposed targeted solutions to mitigate it. We therefore seek effective collaboration that is stage aware and focused, without paying the full cost of direct multi-agent chatting.
To this end, we propose Cochain, a chain-of-collaboration prompting framework for business workflows. Cochain replaces token-intensive direct interaction with retrieval of reusable collaboration artifacts distilled from agents and data (Figure 2). It constructs a collaborative knowledge graph that fuses explicit knowledge from the original datasets with tacit knowledge elicited from specialized agents via counterfactual questioning. To mitigate the influence of irrelevant knowledge Zhou et al. (2024), retrieved evidence is organized into compact causal chains that connect stage-specific information through shared bridge entities, supporting cross-stage reasoning while filtering noise. Cochain also builds a prompts tree that distills reusable, solution-oriented prompts from agent responses and organizes them across workflow stages. Inspired by how the human brain integrates fragmented knowledge into coherent thought Courellis et al. (2024), the prompt tree is designed to provide structured cross-stage guidance: querying the tree yields a prompt chain tailored to the input, guiding the backbone model through stage-specific reasoning in a consistent order. At inference time, Cochain retrieves stage-relevant knowledge, a causal chain from the knowledge graph, and a prompt chain from the prompt tree, then composes them into a structured prompt for the backbone LLM. This produces workflow-aware yet focused answers, balancing insufficient collaboration and excessive collaboration.
@@ 1 Introduction
  • Effective collaboration in business workflows. We identify two opposite collaboration failures in business workflow tasks, insufficient collaboration that misses cross-stage constraints and excessive collaboration that dilutes the core decision with excessive interaction.
  • Cochain framework. We propose Cochain, which enables stage-aware collaboration via reusable artifacts rather than token-heavy interactions. Cochain combines a collaborative knowledge graph with causal-chain retrieval and a prompts tree that composes cross-stage prompt chains for coherent multi-stage reasoning.
  • Empirical evidence and expert support. Experiments across benchmarks and backbones show that Cochain improves quality with efficient inference. Expert evaluation further suggests a smaller model with Cochain can outperform a larger one. We will release our collaborative knowledge graph to support future research.
Figure 2: Overview of Cochain. Offline, we construct a Collaborative Knowledge Graph (CKG) by combining explicit triples from workflow data with tacit triples elicited via counterfactual querying and iterative verification, and build a Prompts Tree through cross-stage prompt distillation. Online, Cochain retrieves stage-relevant knowledge from the CKG, composes causal chains, and leverages the Prompts Tree to propagate necessary constraints across stages, mitigating insufficient collaboration and excessive collaboration.
@@ 2.1 Knowledge Graph Augmented LLMs
LLM reasoning capabilities are critical for high-quality responses Jain et al. (2024). Knowledge graphs, with their structured, explicit, and interpretable nature Zhou et al. (2025), provide clear representations and transparent reasoning paths, guiding LLMs toward deeper reasoning and mitigating limitations Wang et al. (2024). For domain-specific LLMs, knowledge graphs enable continuous knowledge updates to accommodate evolving information Lavrinovics et al. (2025). Applications include MindMap Wen et al. (2024) in medicine and ChatLAW Cui et al. (2024) in law. Business workflows pose an additional challenge of multi-stage, cross-domain knowledge fusion, where irrelevant knowledge can easily overwhelm the core decision and degrade performance Ouyang et al. (2022). To address this, Cochain organizes retrieved evidence into causal chains, improving cross-stage understanding while filtering noise.
@@ 2.2 Prompt Engineering
Prompt engineering guides LLMs to maximize their potential via designed prompts Sahoo et al. (2024). It enables model adaptation to new tasks through in-context learning and instruction following Brown et al. (2020); Li et al. (2023). Well-designed prompts are shown to enhance performance, particularly on complex tasks Wei et al. (2022); Wang et al. (2023); Yao et al. (2023). CoT Wei et al. (2022) enables models to generate intermediate reasoning steps via simple prompts. ToT Yao et al. (2024) explores coherent text units as intermediate problem-solving steps. CPO Zhang et al. (2024a) leverages non-optimal reasoning paths from tree search to speed up inference. However, these methods are typically designed for a single decision context and do not explicitly maintain constraint awareness across stages or support cross-domain coordination in business workflows. Cochain complements this line by distilling cross-stage prompts from agent responses and organizing them into prompts tree for retrieval.